In triangle ABC if segment BC is one inch shoter than segment AB, segment AC is 10 inches shorter than the sum of segment BC and segment AB, and the perimeter of triangle ABC is 72 inches, find the length of segment BC.
A) 20 in.
B) 21 in.
C) 31 in.
D) 32 in.

Answer by fractalier(2101)

Let's call the length of AB, x.
Thus the length of BC is x - 1.
And the length of AC is x + (x - 1) - 10.
Now if we add all the sides we get 72...so
x + (x - 1) + (x + (x - 1) - 10) = 72 and
4x - 12 = 72
4x = 84
x = 21 and x - 1 = 20 is side BC
Choice A.